HelpMePlz:

Find y' if 5y+8x^4=ln(6yx)

Champion:

hint: ln(6yx)=ln6+lny+lnx (multiplication rule of logarithms) and (ln y)'=d(lny)/dy*dy/dx=(1/y)y' (chain rule)

HelpMePlz:

I see, but how do I isolate y? Because it would be 5y-lny=ln6+lnx-8x^4 and I need to find y'

Champion:

You definitely need to isolate y', and y can be used in implicit derivatives unless specifically asked to have the derivative free of y. So don't need to isolate y before deriving. Derive term by term, then isolate y'. The expression will be in terms of x and y and constants.

Champion:

example: xy^2+x^3=0 derive term by term: y^2+2xyy'+3x^2=0 isolate y' 2xyy'=-y^2-3x^2 y'=-(y^2+3x^2)/(2xy)

HelpMePlz:

ohhh ok, I'll try that